# Neural variational inference for cutting feedback during uncertainty propagation

### v1

@@ 4.1 Uniform Kullback-Leibler Approximation Rates of Conditional Normalizing Flows
Table 1 summarizes the tail (large $M$) orders of the spikiness modulus $L_{S}(M)$, the conditional perturbation modulus $L_{CP}(M)$, and the quantile growth $G(M)$ for common conditional families whose parameters (e.g., location, scale, shape) depend continuously on $\eta$.
@@
$\mathrm{GGD}(\mu(\eta),\alpha(\eta),\beta(\eta))$ $O(M^{\max\{0,\beta-1\}})$ $O(M^{\beta}\log M)$ $O(M^{2/\beta})$
@@
Table 1: Tail orders of $L_{S}(M)$, $L_{CP}(M)$, and $G(M)$ for common distribution families. Distribution parameters are $C^{1}$ in $\eta$. Quantile growths are with respect to the Gaussian base.
@@
These examples span a wide range of behavior under the Gaussian-based quantile map $T(\eta,z)=F_{\eta}^{-1}(\Phi(z))$. For a Gaussian conditional law, the base matches the target: $T(\eta,z)=\mu(\eta)+\sigma(\eta)z$ for $|z|\leq M$, so the quantile growth is linear, $G(M)=O(M)$. The spikiness modulus is also linear, $L_{S}(M)=O(M)$ (the score is linear in $\theta$), while the $\eta$–sensitivity can be $O(M^{2})$ because $\log p$ contains a quadratic term in $\theta$. The generalized Gaussian family interpolates smoothly through Gaussian ($\beta=2$), Laplace ($\beta=1$) via the shape parameter $\beta$. The quantiles grow polynomially in the Gaussian base, with order $G(M)=O(M^{2/\beta})$. The log normal and its symmetrized version have tails that decay more slowly than any Gaussian yet faster than any power law. This leads to Gaussian–base quantiles that grow exponentially, summarized by $G(M)=e^{O(M)}$ and is thus an example which sharply meets the exponential quantile bound (16). Thus, families of Gaussian, Laplace, Generalized Gaussian, and log-normal distributions or their finite linear combinations are subsumed in Definition 1.

### v2

@@ Acknowledgements
Table S1 summarizes these three quantities for common conditional families whose parameters (e.g., location, scale, shape) depend continuously on $\eta$.
@@
$\mathrm{GGD}(\mu(\eta),\alpha(\eta),\beta(\eta)\leq 2)$ $O(M^{\max\{0,\bar{\beta}-1\}})$, where $\bar{\beta}=\sup_{\eta}\beta(\eta)$ $O(M^{\bar{\beta}}\log M)$ $O(M^{2/\underline{\beta}})$, where $\underline{\beta}=\inf_{\eta}\beta(\eta)$
@@
Table S1: Tail orders of $L_{S}(M)$, $L_{CP}(M)$, and $G(M)$ for common distribution families. Distribution parameters are $C^{1}$ in $\eta$. Quantile growths are with respect to the Gaussian base.
@@
The rates are derived in Section S3 of the Supplement. These examples span a wide range of behavior under the Gaussian-based quantile map $T(\eta,z)=F_{p}^{-1}(\Phi(z)\,|\,\eta)$. For a Gaussian conditional law, the base matches the target: $T(\eta,z)=\mu(\eta)+\sigma(\eta)z$ for $|z|\leq M$, so the quantile growth is linear, $G(M)=O(M)$. The spikiness modulus is also linear, $L_{S}(M)=O(M)$ (the score is linear in $\theta$), while the $\eta$–sensitivity is $O(M^{2})$ because $\log p$ contains a quadratic term in $\theta$. The generalized Gaussian family interpolates smoothly through Gaussian ($\beta=2$), Laplace ($\beta=1$) via the shape parameter $\beta$. The quantiles grow polynomially in the Gaussian base, with order $G(M)=O(M^{2/\beta})$. The log normal and its symmetrized version have tails that decay more slowly than any Gaussian yet faster than any power law. This leads to Gaussian–base quantiles that grow exponentially, summarized by $G(M)=e^{O(M)}$ and is thus an example which sharply meets the exponential quantile bound (13). Thus, families of Gaussian, Laplace, Generalized Gaussian, and log-normal distributions or their finite linear combinations are subsumed in Assumption 1.